### v1

# Bayesian Concept Bottleneck Models with LLM Priors

Abstract
Concept Bottleneck Models (CBMs) have been proposed as a compromise between white-box and black-box models, aiming to achieve interpretability without sacrificing accuracy. The standard training procedure for CBMs is to predefine a candidate set of human-interpretable concepts, extract their values from the training data, and identify a sparse subset as inputs to a transparent prediction model. However, such approaches are often hampered by the tradeoff between enumerating a sufficiently large set of concepts to include those that are truly relevant versus controlling the cost of obtaining concept extractions. This work investigates a novel approach that sidesteps these challenges: BC-LLM iteratively searches over a potentially infinite set of concepts within a Bayesian framework, in which Large Language Models (LLMs) serve as both a concept extraction mechanism and prior. BC-LLM is broadly applicable and multi-modal. Despite imperfections in LLMs, we prove that BC-LLM can provide rigorous statistical inference and uncertainty quantification. In experiments, it outperforms comparator methods including black-box models, converges more rapidly towards relevant concepts and away from spuriously correlated ones, and is more robust to out-of-distribution samples.
1 Introduction
Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety (Dwork et al.,, 2012; Rudin,, 2018). This has limited their adoption in high-stakes applications such as medicine (Thirunavukarasu et al.,, 2023) and settings where regulatory agencies require algorithms to be explainable (Meskó and Topol,, 2023).
Recent works have explored Concept Bottleneck Models (CBMs) (Koh et al.,, 2020; Kim et al.,, 2023) as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to make predictions. This allows users to understand how exactly each concept contributes to the final prediction and, if necessary, modify the extracted concepts to fix model predictions.
However, there are major limitations in the way CBMs are currently trained. They require human experts to identify a large pool of concepts a priori and annotate the concepts present in each observation. This heavy dependence on human experts is expensive and often impractical. More importantly, it is difficult to ensure this pool of concepts includes the truly relevant concepts. Consequently, resulting CBMs are often outperformed by their black-box counterparts (Yuksekgonul et al.,, 2022) and may even mislead users as to which concepts are truly relevant (Ramaswamy et al.,, 2023).
Consider the task of predicting hospital readmission from patient notes. It is impossible to prespecify all possible related concepts. Not only are there many different concepts, but any single concept can also be refined. For instance, for the concept regarding the “patient’s smoking status,” there is the more general concept regarding “substance use,” the correlated concept of “whether a patient has quit smoking,” and many more. While iterative refinement of candidate concepts may help, it would make learning a CBM even more burdensome.
LLMs have the potential to address these challenges. They are cheaper than human experts, can provide (relatively) high-quality concept annotations, and have sufficient world knowledge to hypothesize useful concepts. However, applying LLMs as a simple stand-in for a human expert does not fully address the challenges of concept discovery. First, LLM queries, though cheaper than human experts, still have a non-zero cost, which practically limits the number of concepts one can explore. Second, searching over concepts using LLMs introduces a “garden of forking paths” and may lead to the selection of spuriously correlated concepts (Gelman and Loken,, 2013). Finally, LLMs are imperfect query engines that can hallucinate, suggest bad candidate concepts, provide incorrect concept annotations, and may not even be self-consistent in their prior beliefs (Wang et al.,, 2023). As such, current methods that rely on LLMs to learn CBMs have struggled to achieve good performance, particularly in areas where the LLM’s prior knowledge is not well-aligned with the truth (Ludan et al.,, 2023; Benara et al.,, 2024).
Figure 1: Overview of Bayesian inference for learning concept bottleneck models with LLM priors (BC-LLM).
To address the aforementioned challenges, we show how wrapping LLMs within a Bayesian framework allows one to iteratively explore concepts in a way that is statistically principled and provides uncertainty quantification. An LLM can serve multiple purposes in this framework: it defines a prior over a potentially infinite set of candidate concepts, proposes concepts to explore during posterior sampling, and annotates observations with the candidate concepts. The procedure, which we refer to as BC-LLM, is designed to be computationally efficient; at each iteration, the LLM proposes a batch of candidate concepts so their respective annotations for each observation can be batched into a single LLM query. We prove that asymptotically, BC-LLM converges to the correct concepts even when the LLM prior is inconsistent or incompatible. Moreover, BC-LLM is highly generalizable and works across multiple modalities (text, images, and tabular data). Experiments across multiple datasets show that BC-LLM outperforms comparator methods and often even black-box models, converges more rapidly towards the true concepts and away from spuriously correlated ones, and is more robust to out-of-distribution (OOD) samples. Finally, when BC-LLM assisted a real-world hospital’s data science team to revise an existing tabular ML model, clinicians found the resulting CBM to be substantially more interpretable and actionable than the prior model.
2 Related work
Prior CBM methods train on observations with hand-annotated (Koh et al.,, 2020; Kim et al.,, 2023; Yuksekgonul et al.,, 2023) and, more recently, LLM-annotated concepts (Patel et al.,, 2023; McInerney et al.,, 2023; Benara et al.,, 2024) for a set of prespecified concepts by human experts. These methods are known to be highly sensitive to the selected set of concepts (Ramaswamy et al.,, 2023). More recent works have suggested using LLMs to prespecify the list of concepts instead (Oikarinen et al.,, 2023; Yang et al.,, 2023), but this requires the LLM to have accurate prior knowledge about the supervised learning task.
Consequently, recent works have hypothesized that taking an iterative approach can more effectively find a set of relevant concepts (Ludan et al.,, 2023; Liu et al., 2024a, ). These methods employ a boosting-like approach, where the LLM is provided a set of misclassified examples and asked to add/revise concepts to improve performance. However, the LLM can typically only review a few observations due to limits on context length, so the resulting concept proposals are far from optimal.
An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts (Kim et al.,, 2018; Yeh et al.,, 2020; Zhao et al.,, 2024; Jiang et al.,, 2023), which can be used to learn an interpretable prediction model. Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model (Doshi-Velez and Kim,, 2017).
Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  (Zhong et al.,, 2023, 2022; Dunlap et al.,, 2024). However, existing methods are primarily suited for simple classification tasks, as they essentially learn a CBM with a single concept rather than a set of concepts.
Finally, recent works have considered combining LLMs with Bayesian techniques (Yang et al.,, 2024; Liu et al., 2024b, ; Liu et al., 2024c, ). Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference (Xie et al.,, 2021; Ye et al.,, 2024), though recent work has proven that ICL is not fully consistent with Bayesian inference (Falck et al.,, 2024). To our knowledge, this work is the first to leverage LLMs in a Bayesian framework for learning CBMs and show that rigorous statistical inference is feasible despite imperfections in LLM reasoning.
3 Method
Consider a dataset $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ are model inputs and $y_{i}$ are labels. Let $\mathbf{X}\coloneq(x_{1},x_{2},\ldots,x_{n})$ and $\mathbf{y}\coloneq(y_{1},\ldots,y_{n})$. A concept $c$ defines a binary- or real-valued function $\phi_{c}$ of a model input, e.g. “Does the doctor note describe an elderly patient?” where $1$=yes and 00=no. A CBM selects a fixed number of $K$ concepts $\vec{c}=(c_{1},\ldots,c_{K})$ and fits a parametric model for $y$ using the extractions $(\phi_{c_{1}}(x),\ldots,\phi_{c_{K}}(x))$. For simplicity of presentation, our discussion will focus on logistic regression (LR)-based CBMs, although other choices are possible. In this case, the data is modeled by
$p\left(Y=1|X=x,\vec{\theta},\vec{c}\right)=\sigma\left(\sum_{k=1}^{K}\theta_{k}\phi_{c_{k}}(x)+\theta_{0}\right)$ (1)
where $\sigma$ is the sigmoid function and $\vec{\theta}=(\theta_{0},\cdots,\theta_{K})$ are the parameters. We refer to $\vec{c}$ as the “support” of the model. Throughout, we use capital and lowercase letters to differentiate between random variables and realized values, respectively. For instance, $C$ refers to a random concept whereas $c$ refers to a particular one.
3.1 Review: Bayesian variable selection
Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated (Ishwaran et al.,, 2008; Castillo et al.,, 2015; Biswas et al.,, 2022). Here, we review the Bayesian solution for this more standard setting. In later sections, we describe extensions to the CBM setting, where there are a potentially infinite number of concepts and a cost associated with extracting each.
[…]
Inference for $p(\vec{c}|\mathbf{y},\mathbf{X})$ is typically achieved through Gibbs sampling. For each epoch, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$. When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs sampling can be used instead (Gilks et al.,, 1995). Rather than replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$. This acceptance probability $\alpha$, delineated in Algorithm 1, depends on the relative posterior probabilities of the candidate and existing concepts and is carefully designed so that the posterior distribution is stationary with respect to the sampling procedure.
Algorithm 1 Metropolis-within-Gibbs
1:Initialize concept set $\vec{c}=(c_{1},\cdots,c_{K})$
2:List of concept sets $L=[]$
3:for $t=1,2,\ldots,T$ do
4:     for $k=1,2,\ldots,K$ do
5:         $c_{k}\leftarrow$ MH-Update($\vec{c}$, $k$) //Update $k$-th concept
6:         Append the current $\vec{c}$ to $L$
7:return $L$
8:function MH-Update($\vec{c}$, $k$)
9:     Propose concept $\check{c}_{k}\sim Q(C_{k};\vec{c}_{-k})$
10:     $\alpha\leftarrow\min\left\{\frac{p((\vec{c}_{-k},\check{c}_{k})|\mathbf{y},\mathbf{X})Q(c_{k};\vec{c}_{-k})}{p(\vec{c}|\mathbf{y},\mathbf{X})Q(\check{c}_{k};\vec{c}_{-k})},1\right\}$
11:     if Accept with probability $\alpha$ then return $\check{c}_{k}$
12:     else return $c_{k}$
3.2 Bayesian variable selection with LLMs
Given the Gibbs procedures described above, there are various ways to integrate LLMs for learning CBMs. To begin, we describe two naïve approaches:
Approach 1: Propose with LLM “prior.” Let the LLM define a conditional prior over concepts, i.e. $p(C_{k}|\vec{c}_{-k})$, and propose concepts from this conditional prior, i.e. $Q(C_{k};\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k})$. As concepts are generated in a data-agnostic fashion, the LLM may produce many irrelevant concepts that have low acceptance probabilities. So while Approach 1 may provide valid inference, it is inefficient and costly.
Approach 2: Propose with LLM “posterior.” Let the LLM conduct posterior inference for the $k$th concept given all the data and all but the $k$th concept through ICL and let the LLM propose from this conditional posterior, i.e. $Q(C_{k};\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$. This simplifies Metropolis-within-Gibbs to Gibbs, which always accepts the proposed concept. As LLMs are not perfect Bayesian inference engines (Falck et al.,, 2024) and may be miscalibrated, this may lead to the inclusion of inappropriate concepts. While efficient, Approach 2 leads to biased sampling.
Instead, we propose to take the middle road, which we argue gives the best of both worlds. At each iteration, we first randomly split the data to construct subset $S$ of size $\lfloor\omega n\rfloor$ and its complement $S^{c}$, for some fraction $\omega$ (e.g. half). Then the LLM is asked to propose a candidate concept based on $\vec{c}_{-k}$ and subset $S$, which can be viewed as approximately sampling from the “partial posterior”
$Q(C_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\approx p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X}).$ (2)
Assuming (2) holds with equality, the Metropolis acceptance probability simplifies to
$\alpha\leftarrow\min\left\{\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})},1\right\}.$
This essentially compares the likelihood of the held-out data with respect to the candidate and existing concepts, much like sample-splitting in frequentist settings. If the CBM with the candidate concept is better, it is accepted; otherwise it is accepted with probability $\alpha$. This procedure, which we refer to as Split-sample Metropolis-within-Gibbs (Algorithm 2) can learn from the data to override erroneous priors and concepts from the LLM, unlike Approach 2. In addition, it is efficient because the LLM will propose relevant concepts based on some of (though not all of) the data, unlike Approach 1.
Algorithm 2 Split-sample Metropolis update
1:function SS-MH-Update($\vec{c}$, $k$)
2:     Sample subset $S$ of size $\lfloor\omega n\rfloor$.
3:     Propose candidate $\check{c}\sim Q(C_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$
4:     $\alpha\leftarrow\min\left\{\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})},1\right\}$
5:     if Accept with probability $\alpha$ then return $\check{c}$
6:     else return $c_{k}$
In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) (Liu et al.,, 2000). Here, the LLM proposes a batch of candidate candidates, and then we sample a candidate from this batch based on their posterior probabilities and consider the sampled candidate for acceptance. This is more cost-effective as it lets us batch concept annotations, as described later.
Algorithm 3 Multiple-try Split-sample Metropolis update
1:function Multi-SS-MH-Update($\vec{c}$, $k$)
2:     Sample subset $S$ of size $\lfloor\omega n\rfloor$.
3:     Propose $\check{c}^{(1)}_{k},...,\check{c}^{(M)}_{k}\sim Q(C_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$
4:     Sample $\check{m}\in\{1,\ldots,M\}$ with probabilities $\propto w_{m}\leftarrow p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}^{(m)}),\mathbf{X})Q(\check{c}^{(m)};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$
5:     $w_{0}\leftarrow p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})Q(c_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$
6:     $\alpha\leftarrow\min\left\{\frac{Q(c_{k};\vec{c}_{-k},\mathbf{y}_{S}\mathbf{X})\sum_{m=1}^{M}w_{m}}{Q(\check{c}_{k}^{(\check{m})};\vec{c}_{-k},\mathbf{y}_{S}\mathbf{X})\sum_{0\leq m\leq M,m\neq\check{m}}w_{m}},1\right\}$
7:     if Accept with probability $\alpha$ then return $\check{c}_{k}^{(\check{m})}$
8:     else return $c_{k}$
Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can show that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition I.1 in the Appendix). More importantly, we can prove that the procedure still converges to the true concepts even in the likely scenario where there is no valid prior.
Theorem 3.1.
Suppose the data is IID. Let $L(\vec{c})\coloneqq\max_{\vec{\theta}}\mathbb{E}\{\log p(Y|X,\vec{\theta},\vec{c})\}$ and $\mathcal{C}^{*}\coloneqq\operatorname{argmax}_{\vec{c}}L(\vec{c}).$ For sample size $n$, let $\Pi_{n}$ denote the set of stationary distributions of the Markov chain defined by running Algorithm 1 with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update. Assuming a Gaussian prior for $\vec{\theta}$ and under regularity conditions (see Assumption H.1 in the Supplement), then $\inf\{\pi(\mathcal{C}^{*})\colon\pi\in\Pi_{n}\}\to 1$ in probability as $n\to\infty$.
3.3 BC-LLM
We now describe the main steps in BC-LLM (Fig 1): the procedure is initialized in Step 0, and each iteration (Steps 1 to 4) samples a CBM. The samples then describe the posterior distribution over CBMs and form an ensemble prediction model (after removing those from the burn-in period). Depending on one’s use case, the search over concepts can be guided by the LLM’s prior knowledge about the target $Y$ by injecting this information into the prompts (e.g. “we are trying to predict readmission…”) or in an “unbiased” fashion by masking the meaning of $Y$ (e.g. “we are trying to predict some label $Y$…”). Example prompts and other implementation details are provided in the Appendix.
Step 0: Initialization and keyphrase extraction
We first use an LLM to extract “keyphrases” describing each observation $x$, which will later be used by the LLM to brainstorm candidate concepts. Here, we use “keyphrases” to refer to short phrases that describe the observation: keyphrases for images describe what’s in the image (e.g. blue eyes), while keyphrases for text data can be phrases that appear in the text or summarize the main ideas (e.g. diabetes) (Papagiannopoulou and Tsoumakas,, 2020). Unlike concepts, keyphrases do not define a function that outputs a value, though they can help one come up with a concept (e.g. “diabetes” is a keyphrase whereas the binary questions “Does the patient currently have diabetes?” or “Does the patient have a family history of diabetes?” are concepts). In addition, this step does not restrict the set of keyphrases one can extract to a prespecified list, which allows for a wide exploration of keyphrases. The tradeoff is that the extracted keyphrases may be, and likely are, an incomplete summary of each observation.
To initialize the CBM sampling procedure, these keyphrases are converted into numeric vectors to fit a (penalized) LR model for the target $Y$, similar to the process for fitting a bag-of-words model. An LLM analyzes the top keyphrases in this LR model to generate an initial set $\vec{c}$ of $K$ concepts. The LLM also extracts these concepts from each observation.
Step 1: Drop a concept & fit a keyphrase model
Each iteration proposes candidate concepts based on some data subset $S$ to replace the $k$th concept for $k\in\{1,\cdots,K\}$. Given the context limits of LLMs, it is unlikely that an LLM can generate a candidate concept from the partial posterior distribution by reviewing each observation in $S$ individually. As such, we will generate a summary by highlighting the top keyphrases in $S$ for predicting $Y$, on top of the existing $K-1$ concepts $\vec{c}_{-k}$.
For continuous outcomes, this can be accomplished by finding the top keyphrases associated with the residuals of a model that predicts $Y$ given $\vec{c}_{-k}$. As the residuals are noisy for binary/categorical outcomes, we fit a (multinomial) LR model of $Y$ instead, where the input features are the keyphrases (again encoded using a bag-of-words approach) and annotated concepts $\vec{c}_{-k}$. The parameters of this “keyphrase model” $f$ are the coefficients for the keyphrases and concepts, denoted $\beta_{W}$ and $\beta_{\vec{c}_{-k}}$, respectively. Given the high-dimensionality of the keyphrases, we penalize $\beta_{W}$ using some penalty function $J$ and estimate $f$ by solving
$\min_{\beta_{\vec{c}_{-k}},\beta_{W}}\frac{1}{n}\sum_{i=1}^{n}\ell\left(y_{i},f\left(x_{i}|\beta_{\vec{c}_{-k}},\beta_{W}\right)\right)+\lambda J(\beta_{W}),$
where $\ell$ is the logistic loss and penalty parameter $\lambda\geq 0$ is selected by cross-validation. We found the ridge penalty (i.e., $J(\beta_{W})=\|\beta_{W}\|_{2}^{2}$) to work well in practice, though other choices for $J$ may be used instead.
Step 2: Query the LLM for candidate concepts Given the summary of the top keyphrases generated from Step 1, the LLM is asked to propose $M$ candidates for the $k$-th concept from the partial posterior $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$. In addition, we ask the LLM to provide the value of the partial posterior probability for each candidate concept. For instance, if the top keyphrases include “chronic kidney disease” and “acute kidney injury,” the LLM may propose a concept like “Does the patient have a history of kidney disease?” with an assigned partial posterior probability of 0.4.
Step 3: Concept annotation Given the batch of candidate concepts, we use the LLM to extract their values for each observation $x$, i.e. $\phi_{c}(x)$ for $c\in\{\check{c}^{(1)}_{k},\ldots,\check{c}^{(M)}_{k}\}$. Because multiple-try Metropolis-within-Gibbs proposes a batch of candidate concepts, we can batch their extractions into a single LLM query for each observation. This makes BC-LLM more cost/query-efficient than ordinary Metropolis-within-Gibbs, which outputs only a single concept per iteration and thus cannot batch extractions.
Step 4: Accept/reject step The final step is to compute the sampling probabilities for the candidate concepts and the acceptance ratio, which involve comparing the proposal probabilities for the existing and candidate concept sets $\vec{c}$ as well as their split-sample posteriors
[…]
To estimate the integral (3), one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops. Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation for (3) (Lewis and Raftery,, 1997), which is discussed in more detail in the Appendix. This essentially approximates the partial and full posterior distributions with Gaussian distributions. As the posterior distribution of LR parameters are asymptotically normal, this approximation becomes increasingly accurate as $n\rightarrow\infty$.
3.4 Computational cost
The key computational cost of BC-LLM is from querying the LLM. Running BC-LLM requires $O(nTK)$ queries, where $T$ is the number of epochs. $K$ is typically small by design, because large $K$ leads to less interpretable CBMs. One may be concerned by $T$, as standard Bayesian inference procedures typically draw thousands of posterior samples. Nevertheless, combined with a greedy warm-start procedure, we found that a small $T$ to be quite effective; our experiments go as low as $T=5$. Generally speaking, if the goal is solely prediction accuracy, a small $T$ may suffice. If the goal is uncertainty quantification for relevant concepts, a larger $T$ (e.g., 10) is helpful for more complete exploration of concepts.
Figure 2: MIMIC evaluation: (a) Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts. Error bars indicate the 95% CI. (b) Dendrograms of concepts learned by BC-LLM with 100 and 800 training observations (left and right, respectively). Labels are shortened concept questions, which generally have the format “Does the note mention the patient having…?” Highlighted labels correspond to the true, color-coded concepts in Section 4.2. Numbers in parentheses describe the proportion of posterior samples with the concept.
4 Experimental results
4.1 Experimental setup
We compare BC-LLM against One-pass summarization, where an LLM performs post-hoc distillation of the keyphrase model into a CBM (Pham et al.,, 2023), and Boosting, where an LLM iteratively adds concepts to a CBM by analyzing observations with large residuals (Ludan et al.,, 2023; Liu et al., 2024a, ). We also compare it to black-box and semi-interpretable models to assess whether BC-LLM can close the performance gap. When the true concepts are known, we also fit an Oracle CBM.
Methods were evaluated with respect to predictive performance as measured by AUC and uncertainty quantification as measured by log likelihood (higher is better). When true concepts are known, we also quantify concept selection rates using “concept precision,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(\hat{C}_{k}\notin\vec{c}^{*}\mid\mathbf{y},\mathbf{X})$, and “concept recall,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(c_{k}^{*}\notin\hat{\vec{C}}\mid\mathbf{y},\mathbf{X})$, where $\hat{\vec{C}}$ is the posterior distribution over concept sets from BC-LLM or a single set of concepts learned by a non-Bayesian procedure. Quantifying a “correct” concept match requires going beyond exact string matching, as there are many possible refinements for a single concept. For instance, for the concept “Does the note mention the patient smoking in the present or the past?”, there are rephrasings (“Does this note mention the patient currently smoking or having smoked?”), polar opposites (“Does the note not mention the currently smoking or having previously smoked?”), or closely overlapping concepts (“Does the note mention the patient currently smokes?”). To quantify if a learned concept was correct, we used the following rule: if the absolute value of the Pearson correlation for LLM annotations of two concepts exceeds 50%, the two concepts are sufficiently similar. We manually confirmed that any pair of learned and true concepts that passed this rule were always closely related.
Experiments used GPT-4o-mini (OpenAI,, 2023) and a protected health information (PHI)-compliant version of GPT-4o for real-world clinical notes in Section 4.4. Additional implementation details and experimental settings are in the Appendix.
4.2 Simulation study: Learning CBMs for clinical notes from MIMIC
We first compared BC-LLM against existing methods in terms of predictive performance and concept selection rates. To assess how well the methods perform even when the LLM has no prior knowledge about the target, we simulated $Y$ using a LR model with five Social Determinants of Health as inputs, which were annotated from real-world patient notes in MIMIC-IV (Johnson et al.,, 2023; Ahsan et al.,, 2021):
1. Does the note mention the patient consuming alcohol in the present or the past?
2. Does the note mention the patient smoking in the present or the past?
3. Does the note mention the patient using recreational drugs in the present or the past?
4. Does the note imply the patient is unemployed?
5. Does the note imply the patient is retired?
CBMs with $K=6$ were trained on 100, 200, 400, and 800 observations and evaluated on 500 held-out observations. This was done for 4 replicates, where the data were randomly selected from the full dataset of 7000 observations. An Oracle CBM was also fit as the concepts are known.
BC-LLM outperforms the comparator methods both in predictive performance and uncertainty quantification (Figure 2a) and, with only 400 observations, performs as well as Oracle. The performance gap between BC-LLM and the comparator methods widens with more training observations because BC-LLM can iteratively refine its concepts. Although Boosting is also an iterative procedure, it struggles to hypothesize relevant concepts because it reviews only a few misclassified observations each iteration and cannot revise concepts added in previous iterations. Additionally, BC-LLM was better at recovering the true concepts compared to existing methods.
Table 1: AUCs of CBMs and black-box models (ResNet) for classifying bird subcategories.
Bird # classes BC-LLM One-pass Boosting ResNet
Albatross 3 0.837 (0.771, 0.894) 0.833 (0.771, 0.890) 0.775 (0.709, 0.842) 0.837 (0.767, 0.897)
Auklet 4 0.884 (0.830, 0.928) 0.865 (0.813, 0.914) 0.733 (0.671, 0.792) 0.831 (0.767, 0.893)
Blackbird 4 0.869 (0.822, 0.912) 0.719 (0.671, 0.771) 0.749 (0.689, 0.801) 0.857 (0.811, 0.900)
Bunting 3 0.932 (0.880, 0.972) 0.629 (0.585, 0.672) 0.681 (0.593, 0.762) 0.919 (0.869, 0.962)
Flycatcher 7 0.862 (0.835, 0.887) 0.800 (0.766, 0.833) 0.784 (0.753, 0.813) 0.817 (0.781, 0.849)
Sparrow 21 0.803 (0.785, 0.819) 0.723 (0.704, 0.739) 0.618 (0.597, 0.640) 0.812 (0.794, 0.830)
Tern 7 0.817 (0.784, 0.849) 0.705 (0.667, 0.740) 0.609 (0.567, 0.651) 0.714 (0.674, 0.751)
The posterior distribution over concepts can be visualized using hierarchical clustering (Figure 2b). Comparing concepts learned given 100 versus 800 observations, we find that the posterior learned by BC-LLM notably contracts towards the true concepts with increasing data.
4.3 Learning to classify birds
To demonstrate the multimodality of BC-LLM, we evaluate how well BC-LLM and other CBM methods can differentiate between bird categories in the CUB-birds image dataset (Welinder et al.,, 2010). [^1] We classify bird categories rather than all subtypes because there is likely a parsimonious CBM for classifying between categories but not between all subtypes. While the exact set of $K$ concepts needed to classify bird categories are unknown, we know certain features that are certainly spurious (e.g. features in the background are irrelevant).
Similar to the prior experiment, we do not tell the LLM that we are predicting bird types. By replacing the bird labels with one-hot encodings, the supervised learning task is more difficult, as the LLM cannot rely on prior knowledge about birds and must conduct an unbiased exploration over concepts. The data is split 50/50 between training and testing (so 30 bird images per category are available for training). Due to cost limitations, seven bird classification tasks were considered. The number of concepts $K$ learned for each task was set to the number of classes, but no smaller than 4 and no greater than 10. For the black-box comparator, we fine-tuned the last layer of ResNet50 pre-trained on ImageNetV2 (He et al.,, 2016; Recht et al.,, 2019).
BC-LLM outperforms all the other CBM learning procedures for every bird type (Table 1). Moreover, it outperforms ResNet50 model in all but one bird. Whereas ResNet50 easily overtrained on the small bird datasets, BC-LLM identified relevant concepts as inputs to a simple parametric model, leading to faster model convergence.
As an example, we show concepts learned by BC-LLM for classifying Bunting birds in Figure 3b. The concepts learned by BC-LLM tended to be more relevant than those from comparator methods. For instance, given all the training data, One-pass summarization checks for a “bird sitting sideways”, “bird on a perch”, and “bird feeding or foraging,” whereas BC-LLM assigned bird position a low posterior weight. Moreover, visualizing the learned concepts helps us understand the learning procedure itself. For instance, we see that BC-LLM learns to downweight spurious features, i.e. concepts relating to backgrounds or bird positions, with accumulating data.
We also assess how robust BC-LLM is in out-of-distribution (OOD) settings, as the Bayesian framing should improve uncertainty quantification. We apply the Bunting CBMs to classify images of the first author’s dog (Fig 3c). Compared to other CBM methods, BC-LLM was more appropriately unsure which Bunting bird to classify the dog as. Interestingly, when the dog is wearing a rainbow unicorn costume (C2), One-pass and Boosting overconfidently classify the dog as the colorful Bunting category (A3) with probabilities greater than 95%. Likewise, when these CBMs were used to classify images of finches (also small colorful birds), the average entropy of BC-LLM was again highest.
Method C1 C2 Finches (e.g., C3)
BC-LLM 0.91 0.90 0.70 (0.53 -0.88)
One-pass 0.06 0.20 0.27 (0.00-0.66)
Boosting 0.00 0.00 0.15 (0.00-0.57)
Figure 3: Bunting bird classification across categories A1-A3. (b) Highlighted labels in the dendrogram correspond to relevant features. Labels are shortened concept questions, generally in the format “Does the image depict…?” Numbers in parentheses describe the proportion of posterior samples with the concept. Left and right dendrograms correspond to learned posteriors for 1/3 and 3/3 of the training data, respectively. (c) To test robustness in OOD settings, the Bunting bird CBMs are asked to classify two images of the first author’s dog and 30 images of finches. Entropies of the model predictions (a measure of model uncertainty) are shown for all the CBM methods. 95% CI is shown for the Finch dataset.
Method AUC (95% CI) Log Lik (95% CI)
BC-LLM 0.64 (0.58, 0.70) -0.45 (-0.50, -0.40)
One-pass 0.59 (0.52, 0.65) -0.46 (-0.51, -0.41)
Boosting 0.59 (0.52, 0.66) -0.46 (-0.51, -0.41)
Bag-of-words 0.52 (0.46, 0.58) -0.47 (-0.52, -0.41)
Figure 4: Learning to augment a readmission risk prediction model for heart failure patients. Labels are shortened versions of concept questions of the format “Does the note mention the patient having…?”
4.4 Augmenting an existing tabular model with real-world clinical notes
Finally, we demonstrate how BC-LLM can help ML developers integrate different data modalities to improve model performance. In particular, the data science team at the Zuckerberg San Francisco General Hospital (ZSFG) has both tabular data and unstructured clinical notes from the electronic health record (EHR). The team has so far fit a model using the tabular data to predict 30-day unplanned readmission risk for patients with congestive heart failure (CHF). Using CBMs, the team wants to assess if clinical notes contain valuable information to extract as additional features. Moreover, the learned model should be interpretable and easy to audit.
To address this, we extend BC-LLM by having the CBM take as inputs (i) the risk prediction from the original tabular model and (ii) $K$ concepts extracted from the clinical notes. This can be viewed as revising the existing model by including additional features. We set to $K=4$. The CBMs were trained on 1000 patients and evaluated on 500 held-out patients. Data in this experiment contained PHI, for which IRB approval was obtained.
The original tabular model, trained on lab and flowsheet values, achieved an AUC of 0.60. BC-LLM revised this model to perform substantially better, achieving an AUC of 0.64 (Figure 4). Moreover, it outperforms comparator methods with respect to both AUC and log likelihood. When the CBMs were shown to four clinicians, [^2] All clinicians work or have worked in ZSFG. Three are in the hospital’s data science team, one of whom (LZ) is a coauthor on this paper. The fourth leads a separate clinical data science research group. They were all interviewed separately. they found that the CBM learned by BC-LLM to be substantially more interpretable and meaningful than the existing tabular model. The clinicians from the hospital’s data science team noted that they had not considered engineering many of the learned concepts on the list. All clinicians noted that many of the learned concepts are known to be causally relevant and would have been difficult to learn from existing tabular data. For example, “substance dependence” is rarely documented directly in the tabular data, and information relevant to determining its value is often missing. Finally, the clinicians highly valued the fact that many of the learned concepts were clinically actionable, such as helping patients better control their diabetes or opting for medication regimes that are easier for patients to follow.
5 Discussion
BC-LLM is a new method for learning CBMs that iteratively proposes concepts using an LLM within a Bayesian framework, which allows for rigorous uncertainty quantification despite LLMs being prone to error and hallucinations. The method is highly general: it is compatible with various data modalities (text, images, and tabular data) and can be extended beyond the settings of binary and multiclass classification. The empirical results show that BC-LLM outperforms existing methods, learns to recover relevant concepts and downweight spurious ones, and can suggest useful features for ML developers to engineer. Code for running BC-LLM and reproducing results in the paper are available at https://github.com/jjfeng/bc-llm.
Acknowledgements
JF was funded through a Patient-Centered Outcomes Research Institute® (PCORI®) Award (ME-2022C1-25619). YT gratefully acknowledges support from NUS Start-up Grant A-8000448-00-00 and MOE AcRF Tier 1 Grant A-8002498-00-00. The authors thank the UCSF AI Tiger Team, Academic Research Services, Research Information Technology, and the Chancellor’s Task Force for Generative AI for their technical support related to the use of Versa API gateway. The authors thank Yifan Mai, Seth Goldman, Andrew Bishara, and Jim Marks for their invaluable feedback on the work. The views presented in this work are solely the responsibility of the author(s) and do not necessarily represent the views of the PCORI®, its Board of Governors or Methodology Committee.
Appendix
Appendix A Multiple-Try Metropolis-Hastings
In this section, we provide intuition for the Multiple-Try Metropolis-Hastings partial posterior method (Algorithm 3). To draw connections with the original Multiple-Try Metropolis-Hastings method (Liu et al.,, 2000), we first work in a completely abstract setting.
Given a state space $\Omega$, stationary distribution $\pi$, and proposal transition kernel $T(y;x)$, define weights $w(x,y)=\pi(x)T(y;x)$. Note that under regular Metropolis-Hastings, the acceptance ratio for proposing the state $y$ given a state $x$ is $\min\{w(y,x)/w(x,y),1\}$.
Classical version.
We briefly describe Liu et al., (2000)’s original method. Suppose the current state is $x$.
  1. 1.
    Draw $z_{1},z_{2},\ldots,z_{M}\sim T(Z;x)$.
  2. 2.
    Sample $y$ from a distribution with probabilities
    $\mathbb{P}\{Y=z_{j}\}=\frac{w(z_{j},x)}{\sum_{i=1}^{M}w(z_{i},x)}.$ (4)
    Let $\check{m}$ denote the sampled index.
  3. 3.
    Sample $z_{1}^{*},\ldots,z_{M-1}^{*}\sim T(Z;y)$, and set $z_{M}^{*}=x$.
  4. 4.
    Accept $y$ with probability
    $\beta(x,y;\mathbf{z},\mathbf{z}^{*})=\min\left\{\frac{\sum_{i=1}^{M}w(z_{i},x)}{\sum_{i=1}^{M}w(z_{i}^{*},y)},1\right\}.$ (5)
One can show that this Markov chain satisfies the detailed balance equations when marginalizing out the intermediate states $z_{1},z_{2},\ldots,z_{M},z_{1}^{*},\ldots,z_{M}$ (see Theorem 1 in Liu et al., (2000)). We would like, however, to minimize the number of proposals necessary, so it is natural to ask whether the algorithm can be modified such that the proposals needed for the backward transitions, $z_{1}^{*},\ldots,z_{M-1}^{*}$, can be omitted. This is indeed possible.
Modified version.
  1. 1.
    Draw $z_{1},z_{2},\ldots,z_{M}\sim T(Z;x)$.
  2. 2.
    Sample $y$ from a distribution with probabilities
    $\mathbb{P}\{Y=z_{j}\}=\frac{w(z_{j},x)}{\sum_{i=1}^{M}w(z_{i},x)}.$ (6)
    Let $\check{m}$ denote the sampled index, $\mathbf{z}=(z_{1},z_{2},\ldots,z_{M})$, $\mathbf{z}^{*}=(z_{1},\ldots,z_{\check{m}-1},x,z_{\check{m}+1},\ldots,z_{M})$, and set $q(y;\mathbf{z})=\mathbb{P}\{Y=y|z_{1},\ldots,z_{M}\}$.
  3. 3.
    Accept $y$ with probability
    $\begin{split}\alpha(x,y;\mathbf{z}_{-\check{m}})&=\min\left\{\frac{p(y\rightarrow\mathbf{z}^{*}\rightarrow x)}{p(x\rightarrow\mathbf{z}\rightarrow y)},1\right\}\\ &=\min\left\{\frac{\pi(y)\prod_{i=1}^{M}T(z_{i};y)q(x;\mathbf{z}^{*})}{\pi(x)\prod_{i=1}^{M}T(z_{i}^{*};x)q(y;\mathbf{z})},1\right\}.\end{split}$ (7)
Proposition A.1.
The modified Multiple-Try MH sampler is reversible.
Proof.
Let $\Phi(y;x)$ denote the actual transition matrix. We want to show
$\pi(x)\Phi(y;x)=\pi(y)\Phi(x;y).$ (8)
To see this, we compute
[…]
Connections between modified and classical versions.
The acceptance probability in the modified version of Multiple-Try MH can be expanded as follows:
$\begin{split}\min\left\{\frac{\pi(y)\prod_{i=1}^{M}T(z_{i};y)q(x;\mathbf{z}^{*})}{\pi(x)\prod_{i=1}^{M}T(z_{i}^{*};x)q(y;\mathbf{z})},1\right\}&=\min\left\{\frac{\pi(y)\prod_{i=1}^{M}T(z_{i};y)\frac{w(x,y)}{\sum_{i=1}^{M}w(z_{i}^{*},y)}}{\pi(x)\prod_{i=1}^{M}T(z_{i}^{*};x)\frac{w(y,x)}{\sum_{i=1}^{M}w(z_{i},x)}},1\right\}\\ &=\min\left\{\frac{\sum_{i\neq\check{m}}w(z_{i},x)+w(y,x)}{\sum_{i\neq\check{m}}w(z_{i},y)+w(x,y)}\cdot\prod_{i=2}^{M}\frac{T(z_{i};y)}{T(z_{i};x)},1\right\}.\end{split}$ (10)
If $T(-,-)$ is invariant in the first argument (as in our application to BC-LLM), then this last formula is exactly equal to $\beta(x,y;\mathbf{z},\mathbf{z}^{*})$, the acceptance probability for the classical version, but if we were to use the same points for both the non-realized forward and backward proposals.
Equivalence with Multi-SS-MH-Update.
Finally, we show that the modified version of Multiple-Try MH described in the previous sections is equivalent to what is implemented in Multi-SS-MH-Update, for a fixed $S$. To see this, we make the replacements:
  • $x\leftarrow\vec{c}$
  • $z_{i}\leftarrow(\check{c}_{k}^{(i)},\vec{c}_{-k})$ for $i=1,\ldots,M$
  • $y\leftarrow(\check{c}_{k}^{(\check{m})},\vec{c}_{-k})$
  • $T(-;-)\leftarrow Q(-;-,\mathbf{y}_{S},\mathbf{X})$
  • $\pi(-)\leftarrow p(-|\mathbf{y},\mathbf{X})$
Plugging these into (10), the acceptance ratio becomes
$\begin{split}&\min\left\{\frac{Q(c_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\sum_{m=1}^{M}p((\check{c}^{(m)},\vec{c}_{-k})|\mathbf{y},\mathbf{X})}{Q(\check{c}_{k}^{(\check{m})};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\sum_{m=1}^{M}p((\check{c}^{(m)},\vec{c}_{-k})|\mathbf{y},\mathbf{X})},1\right\}\\ =~{}&\min\left\{\frac{Q(c_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\sum_{m=1}^{M}p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\check{c}_{k}^{(m)},\vec{c}_{-k}),\mathbf{X})Q(\check{c}_{k}^{(m)};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}{Q(\check{c}_{k}^{(\check{m})};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\sum_{0\leq m\leq M,m\neq\check{m}}p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\check{c}_{k}^{(m)},\vec{c}_{-k}),\mathbf{X})Q(\check{c}_{k}^{(m)};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})},1\right\},\end{split}$ (11)
where the equality comes from (48) and we set $\check{c}_{k}^{(0)}=c_{k}$ for convenience of notation. Observe that this is exactly the formula in Line 6 of Algorithm 3.
Meanwhile, plugging into (6) and using (48), we see that the sampling weights used in the modified version of Multiple-Try MH are equivalent to those in Line 3 of Algorithm 3. This completes the proof of the equivalence.
Appendix B Implementation details for BC-LLM
Here we discuss how the hyperparameters for BC-LLM should be selected:
  • Fraction $\omega$ of data used for partial posterior: Choosing a small $\omega$ may lead to the LLM proposing less relevant concepts, but tends to lead to more diverse proposals. In contrast, a large $\omega$ tends to lead to less diverse proposals because the LLM is encouraged to propose concepts that are relevant to the dataset $\mathcal{D}$, which may not necessarily generalize. In experiments, we found that $\omega=0.5$ provided good results.
  • Number of candidate concepts $M$: More candidate proposals per iteration can allow for more efficient exploration of concepts. The number of candidates is limited by the number of concepts that the LLM can reliably extract in a single batch. There tends to also be diminishing returns, as the first few candidates generated by the LLM based on the top keyphrases tend to be highest quality and most relevant. We found that setting $M=10$ provided good performance.
  • Warm-start and Burn-in: Since Gibbs sampling can be slow to converge, we precede it with a warm-start, which we obtain by updating concepts greedily. That is, we select the concept that maximizes $\operatorname{argmax}p(\gamma|\vec{c}_{-k},\mathbf{y}_{S},\boldsymbol{\Phi})$, instead of sampling from the distribution. In experiments, we run warm-start for one epoch and stored the last 20 iterates as posterior samples; the rest of the samples were treated as burn-in.
  • Number of iterations $T$: Although Monte Carlo procedures for Bayesian inference typically have thousands of samples, this is cost-prohibitive when an LLM must query each observation for a new concept per iteration. In experiments, we found that even setting $T$ as low as $4$ to be quite effective. Generally speaking, if the goal is solely prediction accuracy, a small $T$ may suffice. On the other hand, if the goal is uncertainty quantification for the relevant concepts, one may prefer $T$ on the higher end (e.g. 10) for more complete exploration of concepts.
Appendix C Example prompt for keyphrase extraction (Step 0)
Here is a clinical note: $\langle$ note $\rangle$
Output a list of descriptors that summarizes the patient case (such as aspects on demographics, diagnoses, social determinants of health, etc). For each descriptor, also list as many descriptors that mean the same thing or generalizations of the descriptor. All descriptors, synonyms, and generalizations cannot be more than two words. Output as a JSON in the following format. Do not output more than 30 entries in the JSON:
Appendix D Example prompt for proposing concepts (Step 2)
The goal is to come up with a concept bottleneck model (CBM) that only extracts 5 meta-concepts from patient notes to predict some outcome Y with maximum accuracy. A meta-concept is a binary feature extractor defined by a yes/no question. We have 4 meta-concepts so far:
  1. 1.
    Does the note mention the patient having social support?
  2. 2.
    Does the note mention the patient having stable blood pressure?
  3. 3.
    Does the note mention the patient not experiencing respiratory distress?
  4. 4.
    Does the note mention the patient experiencing substance abuse or dependence?
To come up with the 5th meta-concept, I have done the following: I first fit a CBM on the 4 existing meta-concepts. Then to figure out how to improve this 4-concept CBM, I first asked an LLM to extract a list of concepts that are present in each note, and then fit a linear regression model on the extracted concepts to predict the residuals of the 4-concept CBM. These are the top extracted concepts in the resulting residual model, in descending order of importance:
$\langle$ top words from residual model $\rangle$
To interpret this residual model, a general rule of thumb is that an extracted concept with a large positive coefficient means that the outcome is positively associated with the concept being mentioned in the note, and an extracted concept with a negative coefficient means that the outcome is negatively associated with the concept being mentioned in the note. Remember that notes typically discuss the current state as well as the past medical history of a patient, so a concept being mentioned is in reference to the present OR past state of the patient.
Given the residual model, create cohesive candidates for the 5th meta-concept. Be systematic and consider all the listed concepts in the residual model. Start from the most to the least predictive concept. For each concept, check if it matches an existing meta-concept or create a new candidate meta-concept. Work down the list, iterating through each concept. Clearly state each candidate meta-concept as a yes/no question.
Suggestions for generating candidate meta-concepts: Do not propose meta-concepts that are simply a union of two different concepts (e.g. “Does the note mention this patient experiencing stomach pain or being female?” is not allowed), questions with answers that are almost always a yes (e.g. the answer to “Does the note mention this patient being sick?” is almost always yes), or questions where the yes/no options are not clearly defined (e.g. “Does the note mention this patient experiencing difficulty?” is not clearly defined because difficulty may mean financial difficulty, physical difficulties, etc). Do not propose meta-concepts where you would expect over 95% agreement or disagreement with the 4 existing meta-concepts (e.g. “Does the note mention the patient having high blood pressure?” overlaps too much with “Does the note mention the patient having hypertension?”).
Finally, summarize all the generated candidates for the 5th meta-concept in a JSON. Merge any candidate meta-concepts that are essentially the same (where you would expect over 95% agreement) or essentially opposites (you would expect over 95% disagreement). In the JSON, include a list of comma-separated list of phrases that mean the same or opposite of each candidate meta-concept.
Appendix E Example prompt for annotations for candidate concepts (Step 3)
A simple concept extraction procedure is to obtain binary or categorical concept extractions. In line with findings in Kim et al., (2023), we found that BC-LLM performed better when the LLM quantified its uncertainty when extracting concepts. So if the LLM was unsure about the concept’s value, it was allowed to return a probability. As such, we use prompts like the following:
You will be given a clinical note. I will give you a series of questions. Your task is answer each question with 1 for yes or 0 for no. If the answer to the question is not clearly yes or no, you may answer with the probability that the answer is a yes. Respond with a JSON that includes your answer to all of the questions. Questions:
  1. 1.
    Does the note mention the patient having social support?
  2. 2.
    Does the note mention the patient having stable blood pressure?
  3. 3.
    Does the note mention the patient not experiencing respiratory distress?
  4. 4.
    Does the note mention the patient experiencing substance abuse or dependence?
  5. 5.
    Does the note mention the patient being uninsured?
clinical note: $\langle$ note $\rangle$
Appendix F Experiment details
Note on computation time: When running BC-LLM with GPT-4o-mini-2024-07-18 and GPT-4o, each iteration completed within a few minutes.
F.1 MIMIC
The entire dataset consists of 7043 observations of which we trained on 100, 200, 400, and 800 randomly selected observations and evaluated 500 held-out observations. We used the “Chief Complaint” and “Social History” sections of the notes. The label $Y$ was generated per a LR model, in which
[…]
One-pass summarization and BC-LLM were run to fit CBMs with $K=6$ concepts. BC-LLM was run for $T=5$ epochs. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
F.2 CUB-Birds dataset
To train a bird classifier for $R$ subtypes, One-pass summarization and BC-LLM were run to fit CBMs with $K=\min(10,\max(4,R))$ concepts. BC-LLM was run for a maximum of $T=5$ epochs or 25 iterations, whichever was reached first. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
F.3 Clinical notes and tabular data from ZSFG
To predict readmission risk, the models were trained to analyze the sections “Brief history leading to hospitalization” and the “Summary of hospitalization” in the discharge summary for each patient. One-pass summarization and BC-LLM were run to fit CBMs with $K=4$ concepts. BC-LLM was run for $T=5$ epochs. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration. For the black-box comparator, we used a ResNet50 model pre-trained on ImageNetV2, which ensures ResNet’s training set does not overlap with the CUB-Birds dataset.
For the OOD experiment, the Bunting CBM was applied to 30 images of finches.
Appendix G Laplace Approximation of Split-Sample Posterior
In this section, we describe how to perform Laplace approximation of the split-sample posterior $p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})$ when the likelihood model is logistic and the prior on the coefficient vector $\vec{\theta}$ is a standard normal $\mathcal{N}(0,\gamma^{2}I)$. For simplicity, we omit discussing the constant term $\theta_{0}$. Fixing the list of concepts $\vec{c}$, we let $\Phi$ denote the $n\times K$ matrix whose $(i,j)$-th entry is given by $\Phi_{ij}=\phi_{c_{j}}(x_{i})$. Given any subset of example indices $T\subset[n]$, we can write the joint likelihood for these examples as
$p(\mathbf{y}_{T}|\vec{\theta},\vec{c},\mathbf{X})=\exp\left(\sum_{i\in T}\left(y_{i}\vec{\theta}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}^{T}\Phi_{i\cdot}))\right)\right).$ (12)
Multiplying by the prior for $\vec{\theta}$, the joint conditional distribution for $\mathbf{y}_{T}$ and $\vec{\theta}$ is then:
$p(\vec{\theta},\mathbf{y}_{T}|\vec{c},\mathbf{X})=(2\pi\gamma^{2})^{-1/2}\exp\left(\sum_{i\in T}\left(y_{i}\vec{\theta}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}^{T}\Phi_{i\cdot}))\right)-\frac{\|\vec{\theta}\|_{2}^{2}}{2\gamma^{2}}\right).$ (13)
To form the Laplace approximation to (13), we first denote
$\begin{split}g_{T}(\vec{\theta})&\coloneqq-\log p(\vec{\theta},\mathbf{y}_{S^{c}}|\vec{c},\mathbf{X})\\ &=\sum_{i\in T}\left(-y_{i}\vec{\theta}^{T}\Phi_{i\cdot}+\log(1+\exp(\vec{\theta}^{T}\Phi_{i\cdot}))\right)+\frac{\|\vec{\theta}\|_{2}^{2}}{2\gamma^{2}},\end{split}$ (14)
and its Hessian by
$H_{T}(\vec{\theta})\coloneqq\nabla^{2}g_{T}(\vec{\theta})=\sum_{i\in T}\frac{e^{\vec{\theta}^{T}\Phi_{i\cdot}}}{(1+e^{\vec{\theta}^{T}\Phi_{i\cdot}})^{2}}\Phi_{i\cdot}\Phi_{i\cdot}^{T}+\gamma^{-2}I.$ (15)
Let
$\vec{\theta}_{\operatorname{\operatorname{MAP}},T}\coloneqq\operatorname{argmin}g_{T}(\vec{\theta})$ (16)
be the maximum a posteriori (MAP) estimate. We can then perform a quadratic approximation of the exponent around $\vec{\theta}_{\operatorname{MAP},T}$ to get
$p(\vec{\theta},\mathbf{y}_{T}|\vec{c},\mathbf{X})\approx p(\vec{\theta}_{\operatorname{MAP},T},\mathbf{y}_{T}|\vec{c},\mathbf{X})\exp\left(-\frac{1}{2}(\vec{\theta}-\vec{\theta}_{\operatorname{MAP},T})^{T}H_{T}(\vec{\theta}_{\operatorname{MAP},T})(\vec{\theta}-\vec{\theta}_{\operatorname{MAP},T})\right).$ (17)
Plugging in (13) evaluated at $\vec{\theta}_{\operatorname{MAP},T}$ and integrating out with respect to $\vec{\theta}$ gives
$\begin{split}p(\mathbf{y}_{T}|\vec{c},\mathbf{X})&\approx\exp\left(\sum_{i\in T}\left(y_{i}\vec{\theta}_{\operatorname{MAP},T}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}_{\operatorname{MAP},T}^{T}\Phi_{i\cdot}))\right)+\frac{\|\vec{\theta}_{\operatorname{MAP},T}\|_{2}^{2}}{2\gamma^{2}}\right)\\ &\quad\cdot(2\pi)^{(K-1)/2}\gamma^{-1}\det(H_{T}(\vec{\theta}_{\operatorname{MAP},T}))^{-1/2}.\end{split}$ (18)
We finish by applying the above formulas with respect to $T=[n]$ and $T=S$ and taking ratios:
$\begin{split}&p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})\\ &=\frac{p(\mathbf{y}|\vec{c},\mathbf{X})}{p(\mathbf{y}_{S}|\vec{c},\mathbf{X})}\\ &\approx\exp\left(\sum_{i=1}^{n}\left(y_{i}\vec{\theta}_{\operatorname{MAP}}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}_{\operatorname{MAP}}^{T}\Phi_{i\cdot}))\right)-\sum_{i\in S}\left(y_{i}\vec{\theta}_{\operatorname{MAP},S}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}_{\operatorname{MAP},S}^{T}\Phi_{i\cdot}))\right)\right)\\ &\quad\cdot\exp\left(\frac{\|\vec{\theta}_{\operatorname{MAP},S}\|_{2}^{2}-\|\vec{\theta}_{\operatorname{MAP}}\|_{2}^{2}}{2\gamma^{2}}\right)\left(\frac{\det(H_{S}(\vec{\theta}_{\operatorname{MAP},S}))}{\det(H(\vec{\theta}_{\operatorname{MAP}}))}\right)^{1/2}.\end{split}$ (19)
Remark G.1.
The MAP estimate $\vec{\theta}_{\operatorname{MAP}}$ can be computed by a call to scikit-learn’s LogisticRegression() model, setting C $=2\gamma^{2}$.
Remark G.2.
In the current implementation of our procedure, we use a slightly different method to perform Laplace approximation, which instead approximates the integral in (3).
Appendix H Proof of Theorem 3.1
Recall the definition
$L(\vec{c})\coloneqq\max_{\vec{\theta}}\mathbb{E}_{(X,Y)\sim\nu}\{\log p(Y|X,\vec{\theta},\vec{c})\}.$ (20)
This quantifies how relevant the concepts in $\vec{c}$ are to the prediction of the response $Y$. Note that the definition does not require that the model class is correctly specified (i.e. there does not have to be some $\vec{c}$ and $\vec{\theta}$ so that (1) holds). Before presenting the proof, we first detail the assumptions we make.
Assumption H.1.
  1. (i)
    $Y\in\{0,1\}$ is a binary response;
  2. (ii)
    The data $\mathcal{D}$ is generated i.i.d. from some distribution $\nu$;
  3. (iii)
    The prior on $\vec{\theta}$ is $\mathcal{N}(0,\gamma^{2}I)$;
  4. (iv)
    $\mathcal{C}^{*}\coloneqq\operatorname{argmax}_{\vec{c}}L(\vec{c})$ comprises all permutations of a single vector $\vec{c}^{~{}*}$;
  5. (v)
    There exists some $\Delta>0$ such that $L(\vec{c}^{~{}*})-\operatorname{argmax}_{\vec{c}\notin\mathcal{C}^{*}}L(\vec{c})\geq\Delta$;
  6. (vi)
    Let $\mathcal{C}_{n}\coloneqq\cup_{\vec{c}_{-k},\mathbf{y},\mathbf{X}}\{c\colon Q(c;\vec{c}_{-k},\mathbf{y},\mathbf{X})>0\}$. Assume $\mathcal{C}_{n}$ is a finite set of size at most $\exp(n^{1-\epsilon})$ for some $10<\epsilon<$ 1;
  7. (vii)
    There exists $\eta>0$ such that for each $\vec{c}_{-k}$, there exists $l\in\{1,2,\ldots,K\}$ such that $c_{l}^{*}=\operatorname{argmax}_{c}L((c,\vec{c}_{-k}))$, $c_{l}^{*}\notin\vec{c}_{-k}$, and $Q(c^{*}_{l};\vec{c}_{-k},\mathbf{y},\mathbf{X})\geq\eta$ almost surely for all $\mathbf{y},\mathbf{X}$ with $n$ large enough;
  8. (viii)
    There exists some $B$ such that for any $\vec{c}\in\mathcal{C}_{n}^{K}$, all distinct, we have $\|\vec{\theta}_{\vec{c}}^{*}\|_{2}\leq R$, where
    $\vec{\theta}_{\vec{c}}^{*}\coloneqq\operatorname{argmax}_{\vec{\theta}}\mathbb{E}_{(X,Y)\sim\nu}\{\log p(Y|X,\vec{\theta},\vec{c})\};$ (21)
  9. (ix)
    Let $\Phi_{\vec{c}}=(\phi_{c_{1}}(x),\phi_{c_{2}}(x),\ldots,\phi_{c_{K}}(x))$ for $x\sim\nu$. There exists some $\lambda_{\min}$ such that for any $\vec{c}\in\mathcal{C}_{n,K}$, all eigenvalues of the second moment matrix $\mathbb{E}[\Phi_{\vec{c}}\Phi_{\vec{c}}^{T}]$ are bounded from below by $\lambda_{\min}$. Here, $\mathcal{C}_{n,K}$ denotes the $K$-fold product of $\mathcal{C}_{n}$ but with any vectors containing duplicate concepts removed;
  10. (x)
    We modify the algorithm so that the random subset $S$ can only take one of $\exp(n^{1-\epsilon})$ possibilities $\mathcal{S}_{n}\coloneqq\{S_{1},S_{2},\ldots,S_{B,n}\}$.
Some of these assumptions can be relaxed and are made primarily for the sake of clarity of exposition. First, while we have focused on binary $Y$, the results readily extends to multiclass and continuous $Y$ as well. Second, the Gaussian prior assumption for $\vec{\theta}$ can be relaxed to one that simply assumes some smoothness and coverage properties. Note that for simplicity, we avoid discussing the constant term $\theta_{0}$.
Proof of Theorem 3.1.
We prove the theorem only for SS-MH-Update for simplicity, but it will be clear how to generalize it to Multi-SS-MH-Update.
Step 1: Taking a subsequence. Note that the Markov chain in question corresponds to the list $L$ returned in line 7 of Algorithm 1. Denote this using $(\tilde{\vec{c}}^{~{}t})_{t=1}^{\infty}$. Because of the cyclic nature of the updates, this chain is not time invariant and is hence difficult to work with. We argue that it suffices to study the temporal projection $(\vec{c}^{~{}t})_{t=1}^{\infty}$, where $\vec{c}_{t}=\tilde{\vec{c}}_{tK}$ for $t=0,1,2,\ldots$. This is because every stationary distribution for $(\tilde{\vec{c}}^{~{}t})_{t=1}^{\infty}$ is also stationary for $(\vec{c}^{~{}t})_{t=1}^{\infty}$.
Step 2: Reduction to connected components. Since $(\tilde{\vec{c}}^{~{}t})_{t=1}^{\infty}$ satisfies the detailed balance equations, it is a symmetric Markov chain, i.e. it has an equivalent representation as a flow on a network, in which the vertices are states and the edges correspond to nonzero transition probabilities. This property is inherited by $(\vec{c}^{~{}t})_{t=1}^{\infty}$. Consider a connected component $\mathcal{A}$ of the network. The Markov chain, when restricted to $\mathcal{A}$, is aperiodic, since there must exist a self-loop. Furthermore, it is finite according to Assumption H.1(vi). As such, $(\vec{c}^{~{}t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ (LevinPeresWilmer2006). Indeed, all stationary distributions of $(\vec{c}^{~{}t})_{t=1}^{\infty}$ are convex combinations of such distributions. Hence, it suffices to show the desired property for $\pi_{\mathcal{A}}$.
Step 3: Concentration. In order to analyze $\pi_{\mathcal{A}}$, we need quantitative control over the transition probabilities via concentration. For a given dataset $\mathcal{D}$ of size $n$, we condition on the event guaranteed by Proposition H.3.
Step 4: Convergence. We again focus on a single connected component $\mathcal{A}$ and the goal is to show that with high probability, $\pi_{\mathcal{A}}(\mathcal{C}^{*}\cap\mathcal{A})\geq 1-\delta$ for some $\delta\to 0$ uniformly. We may ignore the rest of the network. Let $a=|\mathcal{A}|$ and $b=|\mathcal{C}^{*}\cap\mathcal{A}|$. Since $\mathcal{A}$ itself is finite, we can enumerate its states starting with those in $\mathcal{C}^{*}\cap\mathcal{A}$. The stationary distribution can be written as a vector $v=(v_{1:b},v_{b+1:a})$, while the Markov chain itself can be represented as a transition matrix $P$, which we write in block form as
$P=\begin{bmatrix}P_{11}&P_{12}\\ P_{21}&P_{22}\end{bmatrix}.$ (22)
The stationarity equation gives us
$v_{b+1:a}(I-P_{22})=v_{1:b}P_{12},$ (23)
from which we obtain
$\|v_{b+1:a}\|_{2}=\left\|v_{1:b}P_{12}(I-P_{22})^{-1}\right\|_{2}\leq\frac{\|P_{12}\|_{op}}{1-\|P_{22}\|_{op}},$ (24)
so long as the denominator on the right hand side is nonzero. To see this, notice that any state $j$ for $b+1\leq j\leq a$ corresponds to a vector $\tilde{\vec{c}}^{~{}(0)}\notin\mathcal{C}^{*}$. However, using Assumption H.1(vii), there is a sequence $(\tilde{\vec{c}}^{~{}(0)},\tilde{\vec{c}}^{~{}(1)},\ldots,\tilde{\vec{c}}^{~{}(K-1)})$ such that for $k=1,\ldots,K$, we have (i) $\tilde{\vec{c}}^{~{}(k)}_{-k}=\tilde{\vec{c}}^{~{}(k-1)}_{-k}$, (ii) $Q(\tilde{c}_{k};\tilde{\vec{c}}^{~{}(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta$, (iii) $L(\tilde{\vec{c}}^{~{}(k)})\geq L(\tilde{\vec{c}}^{~{}(k-1)})$ with equality holding iff $\tilde{\vec{c}}^{~{}(k)}=\tilde{\vec{c}}^{~{}(k-1)}$. Combining (iii) with (32), we see that all acceptance ratios along this sequence are equal to 1, which means that the transition probability from state $j$ to states in $\mathcal{C}^{*}$ is at least
$\prod_{k=1}^{K}Q(\tilde{c}_{k};\tilde{\vec{c}}^{~{}(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta^{K}.$ (25)
The sum of entries in row $j-b$ in $P_{22}$ is the total probability of transition from state $j$ to other states within $\mathcal{A}\backslash\mathcal{C}^{*}$ and hence has value less than $1-\eta^{K}$. This holds for any row. Applying Lemma H.2, we thus get $\|P_{22}\|_{op}\leq 1-\eta^{K}$, which allows us to bound the denominator in (24). Next, every entry in $P_{12}$ is the probability of transition from a state in $\mathcal{C}^{*}$ to a state in $\mathcal{A}\backslash\mathcal{C}^{*}$. Using Step 3 and Assumption H.1(v), the acceptance probability $\alpha$ of such a transition satisfies
$\begin{split}\log\alpha&\leq\max_{\vec{c}\in\mathcal{A}\backslash\mathcal{C}^{*}}\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})-\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c}^{~{}*},\mathbf{X})\\ &\leq-n\Delta+O(n^{1-\epsilon^{\prime}}).\end{split}$ (26)
We thus have
$\|P_{12}\|_{op}\leq\sqrt{(a-b)b}\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}}))\right),$ (27)
which allows us to bound the numerator in (24). Combining this with the earlier bound on the denominator gives
$\|v_{a+1:c}\|_{2}\leq\sqrt{(a-b)b}\eta^{-K}\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}}))\right).$ (28)
We then compute
$\begin{split}\pi_{\mathcal{A}}(\mathcal{C}^{*}\cap\mathcal{A})&=\|v_{1:b}\|_{1}\\ &=1-\|v_{b+1:a}\|_{1}\\ &\geq 1-\sqrt{a-b}\|v_{a+1:c}\|_{2}\\ &\geq 1-\sqrt{b(a-b)^{2}}\eta^{-K}\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}})\right),\end{split}$ (29)
whose error term we can bound as
$\begin{split}\sqrt{b(a-b)^{2}}\eta^{-K}\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}})\right)&\lesssim|\mathcal{C}_{n}|\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}})\right)\\ &\leq\exp(n^{1-\epsilon})\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}})\right)\\ &\leq\exp(O(n^{-\epsilon})).\end{split}$ (30)
Finally, we note that all hidden constants can be chosen to be independent of $n$. ∎
Lemma H.2.
Let $M$ be a $m\times n$ matrix whose entries are non-negative, and suppose $\sum_{j=1}^{n}M_{ij}\leq\beta$ for $i=1,2,\ldots,m$. Then $\|M\|_{op}\leq\beta$.
Proof.
Let $D$ be the $m\times m$ diagonal matrix whose entries are given by $D_{ii}=\sum_{j=1}^{n}M_{ij}$. Then $D^{-1}M$ is a stochastic matrix, which has operator norm at most 1, while $\|D\|_{op}\leq\max_{1\leq i\leq n}\sum_{j=1}^{n}M_{ij}\leq\beta$. Putting these together, we get
$\|M\|_{op}=\|DD^{-1}M\|_{op}\leq\|D\|_{op}\|D^{-1}M\|_{op}\leq\beta,$ (31)
as we wanted. ∎
Proposition H.3 (Concentration of split-sample posterior).
Assuming the conditions in Assumption H.1, then for any $\epsilon^{\prime}<\epsilon/2$, we have
$\sup_{\vec{c}\in\mathcal{C}_{n,K},S\in\mathcal{S}_{n}}\left|\frac{1}{\lceil(1-\omega)n\rceil}\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})-L(\vec{c})\right|=O(n^{-\epsilon^{\prime}})$ (32)
with probability converging to 1.
Proof.
Let us fix $\vec{c}$ for now. We first write
$\begin{split}l_{i}(\vec{\theta})&\coloneqq\log p(Y=y_{i}|X=\mathbf{x}_{i},\vec{\theta},\vec{c})\\ &=y_{i}\vec{\theta}^{T}\Phi_{i\cdot}-\log\left(1+\exp(\vec{\theta}^{T}\Phi_{i\cdot})\right).\end{split}$ (33)
Let $\bar{l}(\vec{\theta})\coloneqq\mathbb{E}_{\mathbf{x}_{i},y_{i}}[l_{i}(\vec{\theta})]$ and set $\vec{\theta}^{*}=\operatorname{argmax}\bar{l}(\vec{\theta})$. For convenience, denote $\Phi=\Phi_{\vec{c}}$. Note that $\Phi$ has entries bounded between 0 and 1 and hence has a second moment matrix with maximum eigenvalue bounded above by $K$. Furthermore, it is a sub-Gaussian random vector with bounded sub-Gaussian norm. Using these properties, we may find a universal value $R^{\prime}$ such that $\mathbb{E}\left[\Phi\Phi^{T}\mathbf{1}(\|\Phi\|_{2}\leq R^{\prime})\right]\succeq\frac{1}{2}\mathbb{E}\left[\Phi\Phi^{T}\right]$. Consider the ball $B(2R)$ of radius $2R$. For any $\vec{\theta}\in B(2R)$, the Hessian of the population log likelihood hence satisfies
$\begin{split}H(\theta)&\coloneqq\nabla^{2}\bar{l}(\vec{\theta})\\ &=-\mathbb{E}\left[\frac{\exp(\Phi^{T}\vec{\theta})}{\left(1+\exp(\Phi^{T}\vec{\theta})\right)^{2}}\Phi\Phi^{T}\right]\\ &\preceq-\mathbb{E}\left[\frac{\exp(\Phi^{T}\vec{\theta})}{\left(1+\exp(\Phi^{T}\vec{\theta})\right)^{2}}\Phi\Phi^{T}\mathbf{1}(\|\Phi\|_{2}\leq M)\right]\\ &\preceq-\frac{1}{2}\inf_{\vec{\theta}\in B(2R),\Phi\in B(M)}\frac{\exp(\Phi^{T}\vec{\theta})}{\left(1+\exp(\Phi^{T}\vec{\theta})\right)^{2}}\mathbb{E}\left[\Phi\Phi^{T}\right]\\ &=\frac{e^{2R^{\prime}R}}{2(1+e^{R^{\prime}R})^{2}}\mathbb{E}\left[\Phi\Phi^{T}\right]\\ &\preceq-\frac{\lambda_{\min}e^{2R^{\prime}R}}{2(1+e^{R^{\prime}R})^{2}}I.\end{split}$ (34)
Using this lower bound on the curvature, we see that
$\bar{l}(\vec{\theta}^{*})-\sup_{\theta\in\partial B(2R)}\bar{l}(\vec{\theta})\geq\frac{\lambda_{\min}R^{2}e^{2R^{\prime}R}}{2(1+e^{R^{\prime}R})^{2}}.$ (35)
Next, we use the fact that $n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)$ is a sub-Gaussian process with increments bounded according to
$\left\|n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)-n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta}^{\prime})-\bar{l}(\vec{\theta}^{\prime})\right)\right\|_{\psi_{2}}\lesssim K\|\vec{\theta}-\vec{\theta}^{\prime}\|_{2}.$ (36)
We can thus use Talagrand’s comparison inequality (vershynin2018high) to get, for any $t>0$, that
$n^{-1/2}\sup_{\vec{\theta}\in B(2R)}\left|\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)\right|\lesssim R\sqrt{K}+Rt$ (37)
with probability at least $1-e^{-t^{2}}$. Take $t=n^{1/2-\epsilon^{\prime}}$ for any $\epsilon^{\prime}<\epsilon/2$. Note that $\hat{l}_{n}(\vec{\theta})\coloneqq n^{-1}\sum_{i=1}^{n}l_{i}(\vec{\theta})$ is concave, with
$\hat{l}_{n}(\vec{\theta}^{*})\geq\bar{l}(\vec{\theta}^{*})-O(n^{-\epsilon^{\prime}}),$ (38)
while
$\begin{split}\sup_{\vec{\theta}\in\partial B(2R)}\hat{l}_{n}(\vec{\theta})&\leq\sup_{\theta\in\partial B(2R)}\bar{l}(\vec{\theta})+O(n^{-\epsilon^{\prime}})\\ &\leq\bar{l}(\vec{\theta}^{*})+\frac{\lambda_{\min}R^{2}e^{2R^{\prime}R}}{2(1+e^{R^{\prime}R})^{2}}+O(n^{-\epsilon^{\prime}}).\end{split}$ (39)
As such, whenever $n$ is large enough, we see that $\hat{l}_{n}$ also attains its global maximum in $B(2R)$ and that this maximum value satisfies
$\left|\sup_{\vec{\theta}\in\mathbb{R}^{K}}\hat{l}_{n}(\vec{\theta})-\bar{l}(\vec{\theta}^{*})\right|=O(n^{-\epsilon^{\prime}}).$ (40)
Recall further that $\bar{l}(\vec{\theta}^{*})=L(\vec{c})$. We next want to integrate out $p(\mathbf{y}|\vec{\theta},\vec{c},\mathbf{X})=\exp(n\hat{l}_{n}(\vec{\theta}))$ with respect to the prior. Using (40), we get the upper bound
[…]
To get a lower bound, we use the fact that the Hessian of the population likelihood also has an lower bound
$H(\theta)\succeq-\|h(\theta)\|_{op}I\succeq-\sqrt{K}I.$ (42)
For any $0<r<R$, this gives a lower bound of
$\inf_{\vec{\theta}\in B(\vec{\theta}^{*},r)}L(\vec{\theta})\geq L(\vec{\theta}^{*})-\sqrt{K}r^{2},$ (43)
where $B(\vec{\theta}^{*},r)$ is the ball of radius $r$ centered at $\vec{\theta}^{*}$. Using (37) again, we get
$\inf_{\vec{\theta}\in B(\vec{\theta}^{*},r)}\hat{l}_{n}(\vec{\theta})\geq L(\vec{\theta}^{*})-\sqrt{K}r^{2}-O(n^{-\epsilon^{\prime}}).$ (44)
The prior density is bounded from below by $\frac{1}{(2\pi\gamma^{2})^{K/2}}\exp(-2R^{2}/\gamma^{2})$ on $B(2R)$. We therefore have
[…]
The maximum is achieved when we set $r=\frac{K^{1/4}}{(2n)^{1/2}}$. Combining (41) and (45) and taking logarithms, we get
$\log p(\mathbf{y}|\vec{c},\mathbf{X})=nL(\vec{c})+O(n^{1-\epsilon^{\prime}}).$ (46)
Performing the same calculation with $\mathbf{y}_{S}$ instead of $\mathbf{y}$ and taking differences, we get
$\left|\frac{1}{\lceil(1-\omega)n\rceil}\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})-L(\vec{c})\right|=O(n^{-\epsilon^{\prime}}).$ (47)
Finally, we take a union bound over all $\vec{c}\in\mathcal{C}_{n,K}$ and $S\in\mathcal{S}_{n}$, noting that all hidden constants can be chosen to be independent of $\vec{c}$. ∎
Appendix I Stationary Distribution Under Consistent Proposals
In this section, we motivate the multiple-try partial posterior Metropolis update (Algorithm 3) in more detail. Notably, we have the following proposition.
Proposition I.1.
Assuming equality holds in (2), the Markov chain defined by running Gibbs sampling (Algorithm 1) with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update has the posterior $p(\vec{C}|\mathbf{y},\mathbf{X})$ as a stationary distribution.
Proof.
We first show this for SS-MH-Update. It suffices to show that the acceptance probability in Line 4 of Algorithm 2 is equal to that of a Metropolis-Hastings filter with $p(\vec{C}|\mathbf{y},\mathbf{X})$ as the target. We compute:
$\begin{split}p((c,\vec{c}_{-k})|\mathbf{X},\mathbf{y})&=\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(c,\vec{c}_{-k}),\mathbf{X})p(\mathbf{y}_{S},(c,\vec{c}_{-k})|\mathbf{X})}{p(\mathbf{y}|\mathbf{X})}\\ &=p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(c,\vec{c}_{-k}),\mathbf{X})p(c|\vec{c}_{-k},\mathbf{X},\mathbf{y}_{S})\frac{p(\vec{c}_{-k},\mathbf{y}_{S}|\mathbf{X})}{p(\mathbf{y}|\mathbf{X})}\end{split}$ (48)
Hence,
$\begin{split}\frac{p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})Q(c;\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}{p((c,\vec{c}_{-k})|\mathbf{y},\mathbf{X})Q(\check{c};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}&=\frac{p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})p((c,\vec{c}_{-k})|\mathbf{X},\mathbf{y})}{p((c,\vec{c}_{-k})|\mathbf{y},\mathbf{X})p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})}\\ &=\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}\end{split}$ (49)
as we wanted.
To show the statement for Multi-SS-MH-Update, we use its equivalence, for a fixed $S$, to the modified Multiple-Try Metropolis-Hastings method described in Appendix A. In Appendix A, we prove that this method satisfies the detailed balance equations for the posterior. Since the detailed balance equations are linear in the transition probabilities, we can marginalize this over $S$ to show that the posterior is stationary with respect for Multi-SS-MH-Update, which draws $S$ uniformly at random. ∎